# Nonreciprocal Photon-Phonon Entanglement in Kerr-Modified Spinning Cavity Magnomechanics

### 09 Dec 2023 v1

@@ Abstract
Cavity magnomechanics has shown great potential in studying macroscopic quantum effects, especially for quantum entanglement, which is a key resource for quantum information science. Here we propose to realize magnons mediated nonreciprocal photon-phonon entanglement with both the magnon Kerr and Sagnac effects in cavity magnomechanics. We find that the mean magnon number can selectively exhibit nonreciprocal linear or nonlinear (bistable) behavior with the strength of the strong driving field on the cavity. Assisted by this driving field, the magnon-phonon coupling is greatly enhanced, leading to the nonreciprocal photon-phonon entanglement via the swapping interaction between the magnons and photons. This nonreciprocal entanglement can be significantly enhanced with the magnon Kerr and Sagnac effects. Given the available parameters, the nonreciprocal photon-phonon entanglement can be preserved at $\sim 3$ K, showing remarkable resilience against the bath temperature. The result reveals that our work holds promise in developing various nonreciprocal devices with both the magnon Kerr and Sagnac effects in cavity magnomechanics.
@@ I Introduction
Magnons Rameshti et al. (2022); Lachance-Quirion et al. (2019); Yuan et al. (2022); Prabhakar and Stancil (2009); Van Kranendonk and Van Vleck (1958), the quanta of collective spin excitations in magnetically ordered materials, especially for the yttrium iron garnet (YIG, $\rm{Y_{3}Fe_{5}O_{12}}$chmidt et al. (2020); Geller and Gilleo (1957); Mallmann et al. (2013), have drawn considerable attention theoretically and experimentally in quantum information science Rameshti et al. (2022); Li et al. (2020). Thanks to high spin density and low collective loss Huebl et al. (2013); Tabuchi et al. (2014); Zhang et al. (2014), magnons in YIG sphere can be strongly coupled to photons in microwave cavities for investigating various phenomena, such as dark modes Zhang et al. (2015); Bi et al. (2019), exceptional point Zhang et al. (2017, 2019a); Zhang and You (2019); Zhao et al. (2020); Cao and Yan (2019); Zhang et al. (2021a); Liu et al. (2019); Sadovnikov et al. (2022); Wang et al. (2023), nearly perfect absorption Rao et al. (2021), unconventional magnon excitations Yuan et al. (2021), , stationary one-way quantum steering Yang et al. (2021a); Guan et al. (2022), and dissipative couplings Wang et al. (2019); Wang and Hu (2020); Harder et al. (2021). With advanced experimental technologies, the magnetostrictive force, overlooked in commonly used dielectric or metallic materials, has been discovered in YIG spheres recently Zhang et al. (2016). This coupling mechanism allows magnons to interact with phonons in vibrated modes of the YIG sphere. Thus, a hybrid cavity magnomechanical (CMM) Zhang et al. (2016) system is built when a YIG sphere meets a cavity. Obviously, this hybrid system combines the individual advantages of magnons, photons and phonons, providing great potential to investigate diverse quantum effects Li et al. (2018, 2019); Hatanaka et al. (2022); Lu et al. (2021); Kani et al. (2022); Huai et al. (2019); Potts (2022). Besides, the magnon Kerr effect (i.e., the Kerr effect of magnons), caused by magnetocrystallographic anisotropy Zhang et al. (2019b); Prabhakar and Stancil (2009), was experimentally demonstrated in cavity magnonics frame Wang et al. (2018, 2016), giving rise to nonlinear cavity magnonics Zheng et al. (2023) as well as Kerr-modified CMM systems Shen et al. (2022). This nonlinearity offers a great power in studying multistability Zhang et al. (2019b); Wang et al. (2018); Shen et al. (2022); Bi et al. (2021); Yang et al. (2021b), long-distance spin-spin interaction Xiong et al. (2022, 2023); Ji and An (2023); Tian et al. (2023), quantum phase transition Liu et al. (2023); Zhang et al. (2021b), and sensitive detection Zhang et al. (2023).
@@
In addition, macroscopic quantum entanglement has garnered significant attention in quantum information science Stannigel et al. (2012); Horodecki et al. (2009); Andersen et al. (2010); Braunstein and van Loock (2005), owing to its wide applications in quantum transduction Tian et al. (2022); Zhong et al. (2022), quantum networking  Cirac et al. (1997); Kimble (2008); Lodahl et al. (2017), quantum sensing  Degen et al. (2017), Bell-state test Marinković et al. (2018); Vivoli et al. (2016), quantum teleportation  Hofer and Hammerer (2015); Hofer et al. (2011), and microwave-optics conversion Xu et al. (2016); Tian and Li (2017); Eshaqi-Sani et al. (2022). To produce such entanglement, nonlinear effects are always required Adesso and Illuminati (2007). This results in the widespread exploration of macroscopic entanglement within nonlinear systems, including nonlinear cavity magnonics Zheng et al. (2023), CMM systems Zuo et al. (2023), and cavity optomechanics Aspelmeyer et al. (2014). Moreover, quantum entanglement can be well protected or enhanced by breaking the Lorentz reciprocity via introducing the Sagnac effect Malykin and Grigorii (2000); Maayani et al. (2018), resulting in the emergence of nonreciprocal entanglement Jiao et al. (2020). Besides the Sagnac effect, the magnon Kerr effect can also be used to achieve nonreciprocal bipartite and tripartite entanglement in cavity optomechanics Chen et al. (2023). This is due to the fact that the magnon Kerr effect can give rise to a positive or negative frequency shift on magnons as well as an additional parametric magnon amplifier under strong driving fields. However, nonreciprocal entanglement with these two nonlinear effects has yet to be revealed to date.
@@
Here we present a scheme to realize a nonreciprocal photon-phonon entanglement in a Kerr-modified spinning CMM system, where the magnon Kerr and the Sagnac effects are both considered. The proposed system consists of the Kerr magnons in the YIG sphere simultaneously coupled to photons in the spinning cavity via the magnetic dipole interaction and phonons in the mechanical mode via magnetostrictive interaction. In this system, the mean magnon number can selectively exhibit linear or nonlinear (bistable) nonreciprocal response under the strong driving field, where the nonreciprocity is induced by the Sagnac effect, and the linear (nonlinear) behavior is caused by the interplay between the magnon Kerr effect and the magnetostrictive interaction. With the enhanced magnetostrictive coupling between the magnons and phonons by the strong driving field, magnons mediated photon-phonon entanglement can be attained via the magnon-phonon and magnon-photon swapping interaction. This achieved entanglement can be nonreciprocally improved by both the magnon Kerr and Sagnac effects with varying the tunable system parameters. Moreover, Such the nonreciprocal entanglement is robust against the bath temperature. With the available parameters, the photon-phonon entanglement can survive at $\sim 3$ K, which is much higher than previous proposals. This work provides opportunities for the development of diverse nonreciprocal devices in Kerr-modified spinning cavity magnomechanics.
@@
This paper is organized as follows. In Sec. II, the model and the system Hamiltonian is described. Then steady-state solution and the effective Hamiltonian are given in Sec. III. In Sec. IV, the nonreciprocal photon-phonon entanglement is studied with system parameters by taking both the Sagnac and magnon Kerr effects into account. Finally, a conclusion is given in Sec. V.
@@
Figure 1: (a) Schematic of the Kerr-modified spinning CMM system. $K_{0}$ is the Kerr coefficient of the magnons, which can be tuned by the direction of the magnetic field $B_{0}$. When the magnetic field is aligned along the crystal axis $[100]~{}([110])$, $K_{0}>0~{}(<0)$. For the spinning cavity, a positive (negative) frequency shift $\Delta_{F}$ is produced bia the Sagnac effect when the driving field is clockwise (counterclockwise). (b) The coupling configuration. The Kerr magnons with the decay rate $\kappa_{m}$ is coupled to both the photons in the spinning cavity with the decay rate $\kappa_{a}$ and the phonons in the mechanical mode with the decay rate $\kappa_{b}$. The corresponding coupling strengths are $g_{ma}$ and $g_{mb}$.
@@ II The model and Hamiltonian
We consider a hybrid Kerr-modified spinning cavity magnomechanical system consisting of a spinning resonator at an angular velocity $\Omega$ holding photons coupled to Kerr magnons in the Kittel mode of a $\mu$m-YIG sphere, where the magnons of the YIG sphere placed in a static magnetic field $B_{0}$ is also coupled to phonons in the mechanical mode [see Fig. 1(a)]. The Hamiltonian of the proposed hybrid system can be written as (setting $\hbar=1$)
@@
Here, $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ ($c$) is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$ in Eq. (3) denotes the relativistic origin of the Sagnac effect, which is small ($\sim 1\%$Malykin and Grigorii (2000); Maayani et al. (2018) and thus can be ignored. The sign ”+()+~{}(-)+ ( - )” in Eq. (3) corresponds to the clockwise (counterclockwise) driving field, where the direction of the cavity spinning is assumed to be along the clockwise direction. This means $\Delta_{F}>0$ ( $\Delta_{F}<0$ ) for the clockwise (counterclockwise) driving field [see Fig.1(a)].
The second term in Eq. (1) related to $K_{0}$ depicts the Kerr nonlinearity of the magnons in the Kittel mode of the YIG sphere, arising from the magnetocrystallographic anisotropy. The Kerr coefficient $K_{0}$ is reversely propotional to the volume of the YIG sphere Zhang et al. (2019b), and it can be tuned either positive or negative by varying the driection of the static magnetic field Wang et al. (2018). Specifically, when the magnetic field is aligned along the crystallographic axis $[100]$ ($[110]$), we have $K_{0}>0~{}(<0)$ Wang et al. (2018). Experimentally, $K_{0}$ can be tuned from $0.05$ nHz to $100$ nHz for the diameter of the YIG sphere from $1$ mm to $100$ ${\rm\mu}$m. The last term in Eq. (1) is the Hamiltonian of the driving field acting on the spinning cavity, where $\epsilon_{d}=\sqrt{2\kappa_{a}P/\omega_{d}}$ is the Rabi frequency, with $P$ being the power and $\omega_{d}$ the frequency. The operators $a~{}(a^{\dagger})$, $b~{}(b^{\dagger})$, and $m~{}(m^{\dagger})$ are the annihilation (creation) operators of the spinning cavity, the mechanical mode, and the Kittel mode, respectively. In the rotating frame respect to $\omega_{d}$, the Hamiltonian in Eq. (1) becomes
@@ III.1 Steady state solution
By defining the frequency detuning of the spinning cavity (Kittel) mode from the driving field, i.e., $\Delta_{a(m)}=\omega_{a(m)}-\omega_{d}$, the dynamics of the considered system with dissipation can be governed by the quantum Langevin equations Walls and Milburn (1998),
@@
By rewriting each operator ($\sigma$) as the sum of its expectation ($\sigma_{s}$) and fluctuation ($\sigma$) in Eq. (III.1), i.e., $\sigma\rightarrow\sigma_{s}+\sigma$, a set of equations related to the operator expectation can be given by
@@
where $\tilde{\Delta}_{m}=\Delta_{m}+2g_{mb}{\rm Re}[b_{s}]$ is the frequency detuning induced by the displacement of the mechanical mode ($2g_{mb}{\rm Re}[b_{s}]$), and $\Delta_{K}=2K_{0}|m_{s}|^{2}$ is the frequency shift caused by the magnon Kerr effect. In the long-time limit, the proposed system reaches its steady state, i.e., $\dot{\sigma}_{s}=0$, so Eq. (III.1) reduces to
@@
Since $\Delta_{F}>0~{}(<0)$ is dependent on the direction of the clockwise (counterclockwise) driving field, so the mean photon number ($|a_{s}|^{2}$) has different values for the opposite driving fields, indicating that $|a_{s}|^{2}$ in the spinning cavity behaves nonreciprocally. This nonreciprocity can indirectly give rise to nonreciprocal mean magnon ($|m_{s}|^{2}$) and phonon ($|b_{s}|^{2}$) numbers beause of the direct coupling between the spinning cavity and the Kittel mode of the YIG sphere, and the magnon-mediated coupling between the spinning cavity and the mechanical mode [see the last two equations in Eq. (III.1)]. It is worth mentioning that such a nonreciprocal situation can also be induced by the magnon Kerr effect, even for a non-spinning cavity ($\Delta_{F}=0$). This is because $K_{0}>0$ or $K_{0}<0$, depending on the direction of the magnetic field, leads to $\Delta_{K}>0$ or $\Delta_{K}<0$. Thus, nonreciprocal mean magnon number is directly obtained [see the last equation Eq. (III.1)] and causes the nonreciprocal mean photon and phonon numbers.

### 12 Apr 2024 v2

@@ Abstract
Cavity magnomechanics has shown great potential in studying macroscopic quantum effects, especially for quantum entanglement, which is a key resource for quantum information science. Here we propose to realize magnon mediated nonreciprocal photon-phonon entanglement, which exhibits asymmetry when opposite magnetic or driving fields are respectively applied to the magnons with the Kerr effect or the photons with the Sagnac effect. We find that the mean magnon number can selectively exhibit nonreciprocal linear or nonlinear (bistable) behavior with the strength of the strong driving field on the cavity. Assisted by this driving field, the magnon-phonon coupling is greatly enhanced, leading to the nonreciprocal photon-phonon entanglement via the swapping interaction between the magnons and photons. This nonreciprocal entanglement can be significantly enhanced with the magnon Kerr and Sagnac effects. Given the available parameters, the nonreciprocal photon-phonon entanglement can be preserved at $\sim 3$ K, showing remarkable resilience against the bath temperature. The result reveals that our paper holds promise in developing various nonreciprocal devices with both the magnon Kerr and Sagnac effects in cavity magnomechanics.
@@ I Introduction
Magnons Rameshti et al. (2022); Lachance-Quirion et al. (2019); Yuan et al. (2022); Prabhakar and Stancil (2009); Van Kranendonk and Van Vleck (1958), the quanta of collective spin excitations in magnetically ordered materials, especially for the yttrium iron garnet (YIG, $\rm{Y_{3}Fe_{5}O_{12}}$chmidt et al. (2020); Geller and Gilleo (1957); Mallmann et al. (2013), have drawn considerable attention theoretically and experimentally in quantum information science Rameshti et al. (2022); Li et al. (2020). Thanks to high spin density and low collective loss Huebl et al. (2013); Tabuchi et al. (2014); Zhang et al. (2014), magnons in a YIG sphere can be strongly coupled to photons in microwave cavities for investigating various phenomena, such as dark modes Zhang et al. (2015); Bi et al. (2019), exceptional points Zhang et al. (2017, 2019a); Zhang and You (2019); Zhao et al. (2020); Cao and Yan (2019); Zhang et al. (2021a); Liu et al. (2019); Sadovnikov et al. (2022); Wang et al. (2023), nearly perfect absorption Rao et al. (2021), unconventional magnon excitations Yuan et al. (2021), stationary one-way quantum steering Yang et al. (2021a); Guan et al. (2022), and dissipative couplings Wang et al. (2019); Wang and Hu (2020); Harder et al. (2021); Yu et al. (2024). With advanced experimental technologies, the magnetostrictive force, originating from the deformation of the sphere’s geometric structure during magnetization Kittel (1958), gives rise to the nonlinear interaction between two modes (magnon mode and phonon mode), which was previously overlooked in commonly used dielectric or metallic materials and has been discovered in YIG spheres recently Zhang et al. (2016). This coupling mechanism allows magnons to interact with phonons in vibrated modes of the YIG sphere. Thus, a hybrid cavity magnomechanical (CMM) Zhang et al. (2016) system is built when a YIG sphere meets a cavity. Obviously, this hybrid system combines the individual advantages of magnons, photons and phonons, providing great potential to investigate diverse quantum effects Li et al. (2018, 2019); Hatanaka et al. (2022); Lu et al. (2021); Kani et al. (2022); Huai et al. (2019); Potts et al. (2021). Additionally, the magnon Kerr effect (i.e., the Kerr effect of magnons), which denotes the nonlinear interaction among the magnon numbers caused by the magnetocrystallographic anisotropy Zhang et al. (2019b); Prabhakar and Stancil (2009), was experimentally demonstrated in the cavity magnonics frame Wang et al. (2018, 2016), giving rise to nonlinear cavity magnonics Zheng et al. (2023) as well as Kerr-modified CMM systems Shen et al. (2022). This nonlinearity offers a great power in studying multistability Zhang et al. (2019b); Wang et al. (2018); Shen et al. (2022); Bi et al. (2021); Yang et al. (2021b), long-distance spin-spin interaction Xiong et al. (2022, 2023); Ji and An (2023); Tian et al. (2023), quantum phase transition Liu et al. (2023); Zhang et al. (2021b), and sensitive detection Zhang et al. (2023).
@@
In addition, macroscopic quantum entanglement has garnered significant attention in quantum information scienceStannigel et al. (2012); Horodecki et al. (2009); Andersen et al. (2010); Braunstein and van Loock (2005); Kounalakis et al. (2022); Zou et al. (2022, 2023); Römling et al. (2023), owing to its wide applications in quantum transduction Tian et al. (2022); Zhong et al. (2022), quantum networking  Cirac et al. (1997); Kimble (2008); Lodahl et al. (2017), quantum sensing  Degen et al. (2017), Bell-state tests Marinković et al. (2018); Vivoli et al. (2016), quantum teleportation  Hofer and Hammerer (2015); Hofer et al. (2011), and microwave-optics conversion Xu et al. (2016); Tian and Li (2017); Eshaqi-Sani et al. (2022). To produce such entanglement, nonlinear effects are always required Adesso and Illuminati (2007). This results in the widespread exploration of macroscopic entanglement within nonlinear systems, including nonlinear cavity magnonics Zheng et al. (2023), CMM systems Zuo et al. (2023), and cavity optomechanics (COM) Aspelmeyer et al. (2014). Moreover, quantum entanglement can be well protected or enhanced in a spinning COM Jiao et al. (2020). This is because the spinning COM allows the emergence of the strong correlation between the photon and the phonon in one chosen direction but weak correlation or complete lack of correlation in the opposite direction. Such direction-dependent entanglement is called nonreciprocal entanglement, which exploits the Sagnac-Fizeau effect to induce an opposite frequency shift on the cavity Malykin and Grigorii (2000); Maayani et al. (2018). Specifically, when the rotation direction of the cavity field differs from the direction of the driving field, the light within the cavity undergoes varying equivalent refractive indices during propagation, resulting in an irreversible refractive index for the clockwise and counterclockwise modes. Correspondingly, the Lorentz reciprocity is broken and nonreciprocal entanglement emerges. Besides the Sagnac effect, the magnon Kerr effect can also be used to achieve nonreciprocal bipartite and tripartite entanglement in cavity optomechanics Chen et al. (2023); that is, by only tuning the direction of the applied magnetic field along the crystallographic axis $[100]$ or $[110]$, asymmetric entanglement is produced. This is due to the fact that the magnon Kerr effect can give rise to a positive or negative frequency shift on magnons as well as an additional parametric magnon amplifier under strong driving fields. However, nonreciprocal entanglement with these two nonlinear effects has yet to be revealed to date.
@@
Here we present a scheme to realize a nonreciprocal photon-phonon entanglement in a Kerr-modified spinning CMM system, where the magnon Kerr and the Sagnac effects are both considered. The proposed system consists of the Kerr magnons in the YIG sphere simultaneously coupled to photons in the spinning cavity via the magnetic dipole interaction and phonons in the mechanical mode via magnetostrictive interaction. In this system, the mean magnon number can selectively exhibit linear or nonlinear (bistable) nonreciprocal response under the strong driving field, where the nonreciprocity is induced by the Sagnac effect, and the linear (nonlinear) behavior is caused by the interplay between the magnon Kerr effect and the magnetostrictive interaction. With the enhanced magnetostrictive coupling between the magnons and phonons by the strong driving field, magnon mediated photon-phonon entanglement can be attained via the magnon-phonon and magnon-photon swapping interaction. This achieved entanglement can be nonreciprocally improved by both the magnon Kerr and Sagnac effects with varying the tunable system parameters. To be more clear, we can interpret this nonreciprocal behavior from the view of our symmetrical operations on the proposed system. Without the magnon Kerr effect (or Sagnac effect), the symmetric photon-phonon entanglement is obtained when opposite magnetic fields along the crystallographic axis of the YIG sphere (or driving fields on the non-spinning cavity) are applied. However, the situation is changed when the magnon Kerr or Sagnac effect is taken into account, that is, the opposite magnetic fields (driving field) give rise to asymmetric entanglement. Since the generated entanglement is dependent on the direction of the magnetic field (driving field) applied to the YIG sphere (cavity), we here clarify that nonreciprocal entanglement is predicted in our proposed system. Moreover, such nonreciprocal entanglement is robust against the bath temperature. With the available parameters, the photon-phonon entanglement can survive at $\sim 3$ K, which is much higher than previous proposals. This paper provides opportunities for the development of diverse nonreciprocal devices in Kerr-modified spinning cavity magnomechanics.
@@
This paper is organized as follows. In Sec. II, the model and the system Hamiltonian are described. Then the steady-state solution and the effective Hamiltonian are given in Sec. III. In Sec. IV, the nonreciprocal photon-phonon entanglement is studied with system parameters by taking both the Sagnac and magnon Kerr effects into account. Finally, a conclusion is given in Sec. V.
@@
Figure 1: (a) Schematic of the Kerr-modified spinning CMM system. $K_{0}$ is the Kerr coefficient of the magnons, which can be tuned by the direction of the magnetic field $B_{0}$. When the magnetic field is aligned along the crystal axis $[100]~{}([110])$, $K_{0}>0~{}(<0)$. For the spinning cavity, a positive (negative) frequency shift $\Delta_{F}$ is produced via the Sagnac effect when the driving field is clockwise (counterclockwise). (b) The coupling configuration. The Kerr magnons with the decay rate $\kappa_{m}$ are coupled to both the photons in the spinning cavity with the decay rate $\kappa_{a}$ and the phonons in the mechanical mode with the decay rate $\kappa_{b}$. The corresponding coupling strengths are $g_{ma}$ and $g_{mb}$.
@@ II The model and Hamiltonian
We consider a hybrid Kerr-modified spinning cavity magnomechanical system consisting of a spinning resonator at an angular velocity $\Omega$ holding photons coupled to Kerr magnons in the Kittel mode of a $\mu$m-YIG sphere, where the magnons of the YIG sphere placed in a static magnetic field $B_{0}$ are also coupled to phonons in the mechanical mode [see Fig. 1(a)]. The Hamiltonian of the proposed hybrid system can be written as (setting $\hbar=1$)
@@
Here, $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ ($c$) is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$ in Eq. (3) denotes the relativistic origin of the Sagnac effect, which is small ($\sim 1\%$Malykin and Grigorii (2000); Maayani et al. (2018) and thus can be ignored. The sign "+"("")"""""+"~{}("-")" + " ( " - " ) in Eq. (3) corresponds to the clockwise (counterclockwise) driving field, where the direction of the cavity spinning is assumed to be along the clockwise direction. This means $\Delta_{F}>0$ ( $\Delta_{F}<0$ ) for the clockwise (counterclockwise) driving field [see Fig.1(a)].
The second term in Eq. (1) related to $K_{0}$ depicts the Kerr nonlinearity of the magnons in the Kittel mode of the YIG sphere, arising from the magnetocrystallographic anisotropy. The Kerr coefficient $K_{0}$ is inversely proportional to the volume of the YIG sphere Zhang et al. (2019b), and it can be tuned either positive or negative by varying the direction of the static magnetic field Wang et al. (2018). Specifically, when the magnetic field is aligned along the crystallographic axis $[100]$ ($[110]$), we have $K_{0}>0~{}(<0)$ Wang et al. (2018). Experimentally, $K_{0}$ can be tuned from $0.05$ to $100$ nHz for the diameter of the YIG sphere from $1$ mm to $100$ ${\rm\mu}$m. The last term in Eq. (1) is the Hamiltonian of the driving field acting on the spinning cavity, where $\epsilon_{d}=\sqrt{2\kappa_{a}P/\omega_{d}}$ is the Rabi frequency, with $P$ being the power and $\omega_{d}$ the frequency. The operators $a~{}(a^{\dagger})$, $b~{}(b^{\dagger})$, and $m~{}(m^{\dagger})$ are the annihilation (creation) operators of the spinning cavity, the mechanical mode, and the Kittel mode, respectively. In the rotating frame with respect to $\omega_{d}$, the Hamiltonian in Eq. (1) becomes
@@ III.1 Steady state solution
By defining the frequency detuning of the spinning cavity (Kittel) mode from the driving field, i.e., $\Delta_{a(m)}=\omega_{a(m)}-\omega_{d}$, the dynamics of the considered system with dissipation can be governed by the quantum Langevin equations Walls and Milburn (1994):
@@
By rewriting each operator ($\sigma$) as the sum of its expectation ($\sigma_{s}$) and fluctuation ($\delta\sigma$) in Eqs. (III.1), i.e., $\sigma\rightarrow\sigma_{s}+\delta\sigma$, a set of equations related to the operator expectation can be given by
@@
where $\tilde{\Delta}_{m}=\Delta_{m}+2g_{mb}{\rm Re}[b_{s}]$ is the frequency detuning induced by the displacement of the mechanical mode ($2g_{mb}{\rm Re}[b_{s}]$), and $\Delta_{K}=2K_{0}|m_{s}|^{2}$ is the frequency shift caused by the magnon Kerr effect. In the long-time limit, the proposed system reaches its steady state, i.e., $\dot{\sigma}_{s}=0$, so Eqs. (III.1) reduce to
@@
Since $\Delta_{F}>0~{}(<0)$ is dependent on the direction of the clockwise (counterclockwise) driving field, so the mean photon number ($|a_{s}|^{2}$) has different values for the opposite driving fields, indicating that $|a_{s}|^{2}$ in the spinning cavity behaves nonreciprocally. This nonreciprocity can indirectly give rise to nonreciprocal mean magnon ($|m_{s}|^{2}$) and phonon ($|b_{s}|^{2}$) numbers because of the direct coupling between the spinning cavity and the Kittel mode of the YIG sphere, and the magnon-mediated coupling between the spinning cavity and the mechanical mode [see the last two equations in Eqs. (III.1)]. It is worth mentioning that such a nonreciprocal situation can also be induced by the magnon Kerr effect, even for a non-spinning cavity ($\Delta_{F}=0$). This is because $K_{0}>0$ or $K_{0}<0$, depending on the direction of the magnetic field, leads to $\Delta_{K}>0$ or $\Delta_{K}<0$. Thus, nonreciprocal mean magnon number is directly obtained [see the last equation in Eqs. (III.1)] and causes the nonreciprocal mean photon and phonon numbers.